OpenStudy (anonymous):

What is the equation of a line that passes through (7,8) and has a slope of -3?

TheSmartOne (thesmartone):

\(\sf\Large y-y_1=m(x-x_1)\) Where \(\sf\Large (x_1,y_1)\) is the point, and \(\Large\sf m\) is the slope

TheSmartOne (thesmartone):

Welcome to OpenStudy! @K.H Can you post the answer choices? :)

OpenStudy (anonymous):

y=-3x+29 y=3x+13y y=1/3x - 29 y=-1/3x - 13

RhondaSommer (rhondasommer):

so for point slope youd have y-8=-3(x-7)

TheSmartOne (thesmartone):

Now you distribute the negative 3. Hint: a(b-c) = ab - ac so what is -3(x-7) = ??

OpenStudy (amity):

Answer is 1st. Option

OpenStudy (amity):

Answer is 1st. Option